OpenStudy (anonymous):

What is the solution to the equation? (-9x+5)^1/2=(3+4x)^1/2

OpenStudy (zehanz):

It is the equation:\[\sqrt{-9x+5}=\sqrt{3+4x}\]When I have to solve this, the first thing I do is not go solving it, but trying to see if there are any restrictions on the variable: 1. -9x+5 must not be negative 2. 3+4x must not be negative If you solve this first, you will see that not every real number is allowed! Now that we have done that, we can go happily squaring both sides to lose the radicals. Then solve -9x+5=3+4x and hope the x you get fits in with rules 1. and 2.!

OpenStudy (anonymous):

2/13, 1/12, -8/5, 13/2 wich of these would be the answer

OpenStudy (zehanz):

If you solve -9x+5=3+4x, you get: 13x=2, so 2/13 seems the right one...

OpenStudy (zehanz):

But you should know how to do that if you're already are solving equations with square roots in them ;)